The speed of a car is 66 km/hr. If each wheel of the car is 140 cm in diameter, find the number of revolutions made by each wheel per minute.
Answer by TimothyLamb(4379)

---
circumference of wheel in meters:
c = 2*pi*r
c = 2*pi*(140/2 cm) * (1/100 m/cm)
c = 4.398229715025710 m
---
by definition:
c = 4.398229715025710 m/rev
---
linear speed of car in meters/minute:
s = 66 km/hr * (1000 m/km) * (1/60 hr/min)
s = 1100 m/min
---
angular speed of wheel:
a = s/c
a = (1100 m/min)/(4.398229715025710 m/rev)
---
answer:
a = 250.1 rev/min
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php